OpenStudy (anonymous):

Find the sum of the geometric sequence. three divided by two, three divided by eight, three divided by thirty two, three divided by one hundred and twenty eight, three divided by five hundred and twelve

OpenStudy (kirbykirby):

The geometric sequence can be represented as: \[ a, ar, ar^2, ar^3,...\] where \(a\) is the first term, and \(r\) is the common ration (this you can find by dividing the 2nd term by the 1st, or the 3rd term by the 2nd.. )

OpenStudy (kirbykirby):

Then you can find the sum of the sequence by using the formula: \[\Large \sum_{x=0}^{n-1}ar^{x}=a\frac{1-r^n}{1-r} \]

OpenStudy (anonymous):

@kirbykirby the choices are one divided by two hundred and fifty six one divided by sixteen one thousand twenty three divided by five hundred and twelve 341

OpenStudy (anonymous):

1/16 1/17 1023/512 341

OpenStudy (kirbykirby):

\[ \large \frac{3}{2}\left( \frac{1-\left(\frac{1}{4}\right)^5}{1-\frac{1}{4}}\right)=?\]